OpenStudy (anonymous):

Write a polynomial function f of least degree that has rational coefficients, a leading coefficient of 1, and the given zeros. 1) -2, -1, 1 2) i, 4 3) i, 2, -√3 4) 1, 4, 1 + √2

OpenStudy (anonymous):

first one is easiest. start with \[p(x)=(x+2)(x+1)(x-1)\] and multiply out

OpenStudy (anonymous):

you should get \[p(x)=x^3+2 x^2-x-2\] for the first one second one you need to know that if \[i\] is a root then so is \[-i\] so again multiply \[p(x)=(x-i)(x+i)(x-4)=(x^2+1)(x-4)\]

OpenStudy (anonymous):

then they get more complicated. third one you need \[p(x)=(x+i)(x-i)(x-2)(x+\sqrt{3})(x-\sqrt{3})\] \[=x^5-2 x^4-2 x^3+4 x^2-3 x+6\]

OpenStudy (anonymous):

last one is... \[p(x)=(x-1)(x-4)(x-(1+\sqrt{2})(x-(1-\sqrt{2}))\]